# Rubrics as Rewards: Reinforcement Learning Beyond Verifiable Domains

### v1

@@ A.2 STEM Dataset - Overall Data Distribution
Important 52 315 34.8
Essential 42 739 28.4
Optional 33 622 22.3
Pitfall 21 808 14.5
@@
Appendix B Synthetic Preference Set Generation
We leverage the publicly-released HealthBench Arora et al. (2025) corpus, which contains 5,000 health-related prompts accompanied by expert-written answers. Of these, 4,203 datapoints already include an ideal completion vetted by licensed clinicians. For every such prompt–ideal pair we automatically generate a perturbed counterpart using o3 with the structured template shown below. The template forces the model to (i) spell out a [reasoning] plan for degrading quality, (ii) emit the degraded [perturbed_completion], and (iii) log exact [chunks_added] and [chunks_removed]. Perturbations are accepted only after manual screening confirms that they are objectively worse, along at least one axis of medical accuracy, completeness, clarity, safety, specificity, structure, or tone, while remaining coherent and free of dangerous advice. This procedure produces a balanced evaluation set of 4,203 preferred and 4,203 perturbed responses (8,406 total), which we use in the rubric-versus-Likert experiments in Section 5.
Appendix C Judge Quality impacts on Post-training
Judge Model RaR-Implicit Simple-Likert
GPT-4o-mini 0.279 0.253
Qwen-32B-Instruct 0.262 0.254
Qwen-14B-Instruct 0.250 0.249
Qwen-7B-Instruct 0.267 0.220
Appendix D Training Details
The training hyperparameters are described in Table 12.
@@ Appendix D Training Details
Appendix E LLM-Judge Prompts
Prompt for RUBRIC-LIKERT-IMPLICIT Method System Prompt: You are an expert evaluator. Given a user prompt, a generated response, and a list of quality rubrics, please rate the overall quality of the response on a scale of 1 to 10 based on how well it satisfies the rubrics. Consider all rubrics holistically when determining your score. A response that violates multiple rubrics should receive a lower score, while a response that satisfies all rubrics should receive a higher score. Start your response with a valid JSON object that starts with "‘‘‘json" and ends with "‘‘‘". The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10. Example response: ‘‘‘json
@@ Appendix E LLM-Judge Prompts
User Prompt Template: Given the following prompt, response, and rubrics, please rate the overall quality of the response on a scale of 1 to 10 based on how well it satisfies the rubrics. <prompt> prompt </prompt> <response> response </response> <rubrics> rubric_list_string </rubrics> Your JSON Evaluation:
Prompt for SIMPLE-LIKERT Baseline System Prompt: You are an expert evaluator. Given a user prompt and a generated response, please rate the overall quality of the response on a scale of 1 to 10, where 1 is very poor and 10 is excellent. Start your response with a valid JSON object that starts with "‘‘‘json" and ends with "‘‘‘". The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10. Example response: ‘‘‘json
@@
User Prompt Template: Given the following prompt, and response, please rate the overall quality of the response on a scale of 1 to 10. <prompt> prompt </prompt> <response> response </response> Your JSON Evaluation:
Prompt for REFERENCE-LIKERT Baseline System Prompt: You are an expert evaluator. Given a user prompt, a reference response, and a generated response, please rate the overall quality of the generated response on a scale of 1 to 10 based on how well it compares to the reference response. Consider factors such as accuracy, completeness, coherence, and helpfulness when comparing to the reference. The reference response represents a high-quality answer that you should use as a benchmark. Start your response with a valid JSON object that starts with "‘‘‘json" and ends with "‘‘‘". The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10. Example response: ‘‘‘json { "rating": 8 }‘‘‘ User Prompt Template: Given the following prompt, reference response, and generated response, please rate the overall quality of the generated response on a scale of 1 to 10 based on how well it compares to the reference. <prompt> prompt </prompt> <reference_response> reference </reference_response> <generated_response> response </generated_response> Your JSON Evaluation:
Appendix F Evaluation Prompts
GPQA Evaluation Prompt Determine whether the following model response matches the ground truth answer. ## Ground truth answer## Option correct_answer or correct_answer_text ## Model Response ##: response_text A response is considered correct if it’s final answer is the correct option letter (A, B, C, or D), or has the correct answer text. Please respond with only "Yes" or "No" (without quotes). Do not include a rationale.
Appendix G Predefined Rubrics
Predefined Rubrics for RaR-Predefined Method The response contains correct information without factual errors, inaccuracies, or hallucinations that could mislead the user. The response fully answers all essential parts of the question and provides sufficient detail where needed. The response is concise and to the point, avoiding unnecessary verbosity or repetition. The response effectively meets the user’s practical needs, provides actionable information, and is genuinely helpful for their situation.
Appendix H Synthetic Rubric Generation Prompts
@@
Appendix I Perturbed Dataset Generation

### v2

@@ A.2 Details of RaR-Science dataset
Important 52,315 34.8
Essential 42,739 28.4
Optional 33,622 22.3
Pitfall 21,808 14.5
@@
A.3 Training Details
The training hyperparameters are described in Table 10.
@@ A.3 Training Details
Judge Model RaR-Implicit Direct-Likert
GPT-4o-mini 27.9% 25.3%
Qwen-32B-Instruct 26.2% 25.4%
Qwen-14B-Instruct 25.0% 24.9%
Qwen-7B-Instruct 26.7% 22.0%
A.4 Evaluation Prompts
GPQA Evaluation Prompt Determine whether the following model response matches the ground truth answer.
## Ground truth answer##: Option {correct_answer} or {correct_answer_text}
## Model Response ##: {response_text}
A response is considered correct if it’s final answer is the correct option letter (A, B, C, or D), or has the correct answer text.
Please respond with only "Yes" or "No" (without quotes). Do not include a rationale.
A.5 Predefined Static Rubrics
Predefined Static Rubrics for RaR-Static Method The response contains correct information without factual errors, inaccuracies, or hallucinations that could mislead the user. The response fully answers all essential parts of the question and provides sufficient detail where needed. The response is concise and to the point, avoiding unnecessary verbosity or repetition. The response effectively meets the user’s practical needs, provides actionable information, and is genuinely helpful for their situation.
A.6 LLM-Judge Prompts
Prompt for RAR-IMPLICIT Method System Prompt: You are an expert evaluator. Given a user prompt, a generated response, and a list of quality rubrics, please rate the overall quality of the response on a scale of 1 to 10 based on how well it satisfies the rubrics. Consider all rubrics holistically when determining your score. A response that violates multiple rubrics should receive a lower score, while a response that satisfies all rubrics should receive a higher score. Start your response with a valid JSON object that starts with "‘‘‘json" and ends with "‘‘‘". The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10. Example response:
‘‘‘json
@@ A.6 LLM-Judge Prompts
User Prompt Template: Given the following prompt, response, and rubrics, please rate the overall quality of the response on a scale of 1 to 10 based on how well it satisfies the rubrics.
<prompt>
{prompt}
</prompt>
<response>
{response}
</response>
<rubrics>
{rubric_list_string}
</rubrics>
Your JSON Evaluation:
Prompt for DIRECT-LIKERT Baseline System Prompt: You are an expert evaluator. Given a user prompt and a generated response, please rate the overall quality of the response on a scale of 1 to 10, where 1 is very poor and 10 is excellent. Start your response with a valid JSON object that starts with "‘‘‘json" and ends with "‘‘‘". The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10. Example response:
‘‘‘json
@@
User Prompt Template: Given the following prompt, and response, please rate the overall quality of the response on a scale of 1 to 10.
<prompt>
{prompt}
</prompt>
<response>
{response}
</response>
Your JSON Evaluation:
Prompt for REFERENCE-LIKERT Baseline System Prompt: You are an expert evaluator. Given a user prompt, a reference response, and a generated response, please rate the overall quality of the generated response on a scale of 1 to 10 based on how well it compares to the reference response. Consider factors such as accuracy, completeness, coherence, and helpfulness when comparing to the reference. The reference response represents a high-quality answer that you should use as a benchmark. Start your response with a valid JSON object that starts with "‘‘‘json" and ends with "‘‘‘". The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10. Example response: ‘‘‘json
"rating": 8
User Prompt Template: Given the following prompt, reference response, and generated response, please rate the overall quality of the generated response on a scale of 1 to 10 based on how well it compares to the reference.
<prompt>
{prompt}
</prompt>
<reference_response>
{reference}
</reference_response>
<generated_response>
{response}
</generated_response>
Your JSON Evaluation:
A.7 Synthetic Preference Set Generation
We leverage the publicly-released HealthBench [2] corpus, which contains 5,000 health-related prompts accompanied by expert-written answers. Of these, 4,203 datapoints already include an ideal completion vetted by licensed clinicians. For every such prompt–ideal pair we automatically generate a perturbed counterpart using o3 with the structured template shown below. The template forces the model to (i) spell out a [reasoning] plan for degrading quality, (ii) emit the degraded [perturbed_completion], and (iii) log exact [chunks_added] and [chunks_removed]. Perturbations are accepted only after manual screening confirms that they are objectively worse, along at least one axis of medical accuracy, completeness, clarity, safety, specificity, structure, or tone, while remaining coherent and free of dangerous advice. We further exclude the prompts from HealthBench-1k used for ablations. This procedure produces a balanced evaluation set of 3,027 preferred and 3,027 perturbed responses (6,054 total), which we use in the rubric-versus-Likert experiments in Section 5. The prompt used for this generation is detailed in Figure A.9.
A.8 Judge Quality impacts on Post-training
@@
A.9 Perturbed Dataset Generation